A sample of gas of mass 1.81g occupies a volume of 629mL at 23°C and 1 atm pressure. What is the molar mass of the gas? (R 0.0821 L. atm/mol-K)
Answer #1

Given:

P = 1.0 atm

V = 629.0 mL

= (629.0/1000) L

= 0.629 L

T = 23.0 oC

= (23.0+273) K

= 296 K

find number of moles using:

P * V = n*R*T

1 atm * 0.629 L = n * 0.0821 atm.L/mol.K * 296 K

n = 2.59*10^-2 mol

mass(solute)= 1.81 g

use:

number of mol = mass / molar mass

2.59*10^-2 mol = (1.81 g)/molar mass

molar mass = 69.9 g/mol

Answer: 69.9 g/mol